Merone Company allocates materials handling cost to the company's two products using the below data:

Modular Homes Prefab Barns
Total expected units produced 6,200 9,200
Total expected material moves 620 220
Expected direct labor-hours per unit 820 320

The total materials handling cost for the year is expected to be $321,120.

If the materials handling cost is allocated on the basis of material moves, the total materials handling cost allocated to the modular homes is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $237,019.80

  • $230,981.05

  • $220,770.00

  • $164,265.00
Answer #1

Total material moves = 620 + 220 = 840

Total material handling cost allocated to Modular Homes = $321,120 X (620 / 840) = $237,019.80

1st option